OpenStudy (anonymous):

PLEASE HELP!!!! Let f(x) = square root of x, g(x) = x + 9 and h(x) = 4/x. Find the domain of the following function then simplify the expression. h/g(x)

OpenStudy (anonymous):

h/g(x) = (4/x)/(x+9) = 4/(x(x+9)) this means that x can be anything except for 0, -9, this can be written as R-{0}{-9} (all real numbers except for 0,-9

OpenStudy (anonymous):

Thanks!